Two identical metallic spheres are suspended as bobs of a simple pendulum. Each of the bobs has a charge of  If each string is 1m long and makes an angle of 12 degrees to the vertical, (a) what is the electric force between the spheres? (b) The bobs touched each other and then separated. What then is the electric force between the sphere?


1
Expert's answer
2022-02-15T11:02:11-0500

We know that

"q=q'=7.5\times10^{-9}c"

l=1m

"\theta=12°"

"F=\frac{kq_1q_2}{r^2}"

"r=2lsin\theta"


"F=\frac{9\times10^9\times(7.5\times10^{-9})^2}{(2\times1sin12°)^2}=2.92\times10^{-6}N"

(B)

"F'=\frac{kq_1q_2}{r^2}"


"F'=\frac{9\times10^9\times(7.5\times10^{-9})^2}{(2\times1sin12°)^2}=2.92\times10^{-6}N"

"F=F'"

Force not change